# Cost-Efficient Serving of LLM Agents via Test-Time Plan Caching

### v1

@@ A.1 Algorithms for Framework Design
1:Query $q$, Context $ctx$, Cache $C$
@@
3:$keyword\leftarrow\text{ExtractKeyword}(q)$triangleright Extract keyword using a small LM
@@
5:     $o,C\leftarrow\text{HandleCacheHit}(q,ctx,C[keyword],C)$triangleright Algorithm 2
6$\triangleright$ Cache miss (Figure 2(b))
7:     $o,C^{\prime}\leftarrow\text{HandleCacheMiss}(q,ctx,keyword,C)$triangleright Algorithm 3
@@
9:return $o,C^{\prime}$triangleright Return response and possibly updated cache
@@
1:Query $q$, Context $ctx$, Plan Template $template$, Plan Cache $C$
@@
3:$responses\leftarrow\emptyset$triangleright Initialize actor LM response to be empty
4:$plan,o\leftarrow\text{LightLM}(q,template,responses)$triangleright Adapt the retrieved template to be a task-specific plan using a lightweight model
@@
7:     $response\leftarrow\text{ActorLM}(q,ctx,plan)$triangleright Execute the plan based on context
@@
9:     $plan,o\leftarrow\text{LightLM}(q,template,responses)$triangleright Generate the final output or a new adapted plan
@@
1:Query $q$, Context $ctx$, Plan Template $template$, Plan Cache $C$
@@
3:$log\leftarrow\emptyset$triangleright Initialize the execution log to be empty
4:$responses\leftarrow\emptyset$triangleright Initialize actor LM response to be empty
5:$plan,o\leftarrow\text{PlannerLM}(q,responses)$triangleright Generate initial plan with full model
@@
8:     $response\leftarrow\text{ActorLM}(q,ctx,plan)$triangleright Execute the plan based on context
@@
10:     $log\leftarrow log\cup\{(plan,ctx,response)\}$triangleright Update the log
11:     $plan,o\leftarrow\text{PlannerLM}(q,responses)$triangleright Generate the final output or a new plan
@@
13:$log\leftarrow log\cup\{o\}$triangleright Update the log
14:$template\leftarrow\text{GenerateTemplate}(log,keyword)$triangleright Create reusable plan template based on execution log
15:$C^{\prime}\leftarrow C$
16:$C^{\prime}[keyword]\leftarrow template$triangleright Store template in cache
@@ B.4.1 Agent Prompts
We use the same prompts from the Minion protocol in the Minions project [35].

### v2

@@ A.1 Algorithms for Framework Design
1:Query $q$, Context ctxctx, Cache $C$
@@
[…]
@@
5:  $o,C\leftarrow\text{HandleCacheHit}(q,ctx,C[keyword],C)$ $\triangleright$ Algorithm 2
[…]
7:  $o,C^{\prime}\leftarrow\text{HandleCacheMiss}(q,ctx,keyword,C)$ $\triangleright$ Algorithm 3
@@
9:return $o,C^{\prime}$ $\triangleright$ Return response and possibly updated cache
@@
1:Query $q$, Context ctxctx, Plan Template templatetemplate, Plan Cache $C$
@@
[…]
@@
7:  $response\leftarrow\text{ActorLM}(q,ctx,plan)$ $\triangleright$ Execute the plan based on context
@@
9:  $plan,o\leftarrow\text{LightLM}(q,template,responses)$ $\triangleright$ Generate the final output or a new adapted plan
@@
1:Query $q$, Context ctxctx, Plan Template templatetemplate, Plan Cache $C$
@@
[…]
@@
8:  $response\leftarrow\text{ActorLM}(q,ctx,plan)$ $\triangleright$ Execute the plan based on context
@@
10:  $log\leftarrow log\cup\{(plan,ctx,response)\}$ $\triangleright$ Update the log
11:  $plan,o\leftarrow\text{PlannerLM}(q,responses)$ $\triangleright$ Generate the final output or a new plan
@@
[…]
@@ B.4.1 Agent Prompts
We use the same prompts from the Minion protocol in the Minions project [narayan2025minions].
@@ 3. Plan (round 2): The working capital ratio can be determined by:
Appendix E Extended Discussion of Limitations, Societal Impact, and Future Directions
Limitations and Challenges
Broader Impact and Societal Implications
Future Directions